### v1

# IAP: Improving Continual Learning of Vision-Language Models via Instance-Aware Prompting

Abstract
Recent pre-trained vision-language models (PT-VLMs) often face a Multi-Domain Class-Incremental Learning (MCIL) scenario in practice, where several classes and domains of multi-modal tasks are incrementally arrived. Without access to previously learned tasks and unseen tasks, memory-constrained MCIL suffers from forward and backward forgetting. To alleviate the above challenges, parameter-efficient fine-tuning techniques (PEFT), such as prompt tuning, are employed to adapt the PT-VLM to the diverse incrementally learned tasks. To achieve effective new task adaptation, existing methods only consider the effect of PEFT strategy selection, but neglect the influence of PEFT parameter setting (e.g., prompting). In this paper, we tackle the challenge of optimizing prompt designs for diverse tasks in MCIL and propose an Instance-Aware Prompting (IAP) framework. Specifically, our Instance-Aware Gated Prompting (IA-GP) module enhances adaptation to new tasks while mitigating forgetting by dynamically assigning prompts across transformer layers at the instance level. Our Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP) improves the task adaptation process by determining an accurate task-label-related confidence score for each instance. Experimental evaluations across 11 datasets, using three performance metrics, demonstrate the effectiveness of our proposed method. Code can be found at https://github.com/FerdinandZJU/IAP.
1 Introduction
Recent years have witnessed a great development of deep neural networks in numerous multi-modal applications, where all the required data are available simultaneously for training on various tasks [52]. Nevertheless, real-world applications usually meet a challenging Multi-Domain Class-Incremental Learning (MCIL) scenario [48]: 1) tasks of multi-modal data are acquired incrementally, 2) the task distribution and target classes vary across tasks. Since the learned tasks data are un-available in a memory-constrained MCIL, multi-modal models suffer from the catastrophic forgetting phenomenon [52] from two aspects: backward forgetting and forward forgetting [51], as shown in Figure 1.
Figure 1: Illustration of backward forgetting and forward forgetting. During the current learning process (e.g., Flowers), backward forgetting in traditional incremental learning refers to the degradation of seen tasks (e.g., Aircraft and DTD). In contrast, forward forgetting manifests as a decline in zero-shot generalization performance on unseen tasks (e.g., Food and SUN397).
Due to the remarkable zero-shot ability of vision-language models (VLMs) [12] (e.g., CLIP [33]), existing works have explored the multi-modal class incremental learning with pre-trained VLMs and parameter-efficient fine-tuning (PEFT) techniques (e.g., prompt, adapter, LoRA) [52]. Typically, these pre-trained model-based multi-modal class incremental learning works usually maintain a fixed pre-trained VLM and incrementally learn few task-specific parameters to adapt to various encountered tasks. Since the memory of MCIL is constrained, such pre-trained model-based methods mainly focus on designing an appropriate PEFT strategy for effective adaptation on various tasks. Both the type of PEFT strategy (prompt, adapter or LoRA) and the parameter setting of PEFT (where and how many prompts/adapters/LoRA to integrate) influence the performance of learning a new task in MCIL.
Existing MCIL works have widely discussed the effect of PEFT strategy selection on MCIL performance, but overlook the analysis of the PEFT parameter settings. These works typically employ a manually designed PEFT parameter setting and apply them consistently across the incremental learning of different tasks. For instance, when learning a new task, prompt-based MCIL methods (e.g., DIKI [36]) consistently use the same number of prompts at the same positions within the pre-trained VLM (an operation we refer to as prompting). However, our observations suggest that the optimal prompting strategy for adapting to different tasks varies within the MCIL framework. Thus, the key challenge lies in developing a flexible and adaptive prompting strategy to enhance task adaptation.
Motivated by the above observations, we propose a novel Instance-Aware Prompting (IAP) framework for MCIL, which dynamically allocates prompt positions and weights at the instance level. Specifically, we introduce an Instance-Aware Gated Prompting (IA-GP) strategy that determines whether to apply prompts in each transformer layer based on instance-specific features. Furthermore, we propose an Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP) strategy to compute more reliable confidence scores at the instance level, which are then used as prompt weights in subsequent operations. By integrating above two modules, we solve the problems of fixed prompting strategy of conducting incremental learning on pre-trained VLMs.
The contributions of our proposed IAP approach can be summarized in threefold:
  • We design an Instance-Aware Gated Prompting strategy to address the challenge of determining where to prompt. We enhance PEFT techniques, enabling the model to dynamically allocate prompting positions at the instance level, thereby improving the models incremental learning capability.
  • We introduce Instance-Aware Class-Distribution-Driven Prompting to derive more reliable prompt weights. To enhance model performance across diverse instances, we employ a two-stage distribution modeling approach that operates at both the task and class levels during inference.
  • Extensive experiments on benchmark datasets demonstrate that our IAP method achieves state-of-the-art performance while utilizing only 1.18% of the training parameters, outperforming all existing approaches.
2 Related Works
Incremental learning. Conventional incremental learning can be categorized into three types: task-incremental learning (TIL) [17, 38, 51], domain-incremental learning (DIL) [17, 38, 50], and class-incremental learning (CIL) [17, 38, 49, 11]. Among these, task-incremental learning requires explicit task identifies during inference, while class-incremental learning requires non-overlapping classes across different tasks. There are three technologies in incremental learning. The first is regularization-based methods, such as Elastic Weight Consolidation (EWC) [20] and Memory Aware Synapses (MAS) [1]. These methods introduce regularization terms to constrain optimization directions and mitigate catastrophic forgetting. The second one is rehearsal-based methods [26, 34, 43, 15, 24, 16, 30], including Learning without Forgetting (LwF) [26] and Incremental Classifier and Representation Learning (iCaRL) [34]. These approaches generally require additional storage to retain either previous model parameters or representative samples from seen tasks. The third one is network expansion-based methods [29, 46, 45, 25], with Dynamic Expandable Networks (DEN) [46] being a representative approach. These methods dynamically expand the neural network structure to accommodate new tasks. Recently, observing the strong generalization of pre-trained models, some incremental learning approaches are designed to fine-tune the pre-trained models [19, 37, 53, 4, 13], which called parameter-efficient fine-tuning (PEFT). PEFT techniques achieve adaptation for downstream tasks by fine-tuning only a small number of model parameters or maintaining a limited set of additional parameters while keeping the majority of the pre-trained model’s parameters frozen. PEFT technology significantly reduces computational and storage costs and mitigates the catastrophic forgetting. Notable related methods include L2P [41], DualPrompt [40], S-Prompt [39], and CODA-Prompt [35]. However, these methods only consider stable distributions and lack of the capability to learn tasks with distinct distributions. Unlike conventional approaches, we focus on a novel benchmark called Multi-Domain Class-Incremental Learning (MCIL), where data distributions are distinct and arrive continually.
Downstream tasks of vision-language models. Visual Language Models (VLMs) [18, 32] have achieved remarkable progress in multimodal research, successfully enabling cross-modal understanding and generation capabilities through joint modeling of visual inputs and natural language. With the explosive growth of online data, recent years have witnessed further advancements in VLMs driven by larger-scale models and more extensive datasets. However, when researchers attempt to fine-tune the VLMs for a downstream task, an inevitable degradation occurs in their zero-shot capabilities of other tasks [23, 5]. Various approaches have been proposed to mitigate the degradation. ZSCL [51] leverages the knowledge distillation technology [26], treating the original VLM as a teacher model and distilling knowledge into the fine-tuned model through constructed “wild” [3] datasets (e.g., ImageNet [7]). Although ZSCL partially alleviates forgetting of pre-trained knowledge, it requires additional storage space for “wild” datasets, and its full-parameter fine-tuning strategy incurs substantial computational costs. Alternative approaches such as MoE-Adapter [47] and DIKI [36] employ PEFT technology, updating only a small set of parameters to reduce the forgetting of pre-trained knowledge. Nevertheless, these methods continue to encounter challenges associated with computational resource overhead and the under-exploration of instance-aware features. Our proposed approach overcomes these limitations by introducing an instance-aware prompting strategy, significantly enhancing model performance.
Figure 2: Illustration of Instance-Aware Gated Prompting strategy. The figure illustrates the processing workflow of the IA-GP strategy applied to an instance ”abbey”. Initially, the image is segmented into patches, which are then fed into the visual encoder of the CLIP model. Simultaneously, corresponding textual category descriptions are processed by the textual encoder. The IA-GP strategy utilizes trainable prompt pools for both visual and textual modalities, while prompts from previously seen tasks are kept frozen. The strategy leverages the visual features from the original CLIP encoders, denoted as $f_{v}=F_{v}(x)$, as input of prompt gate module. Hard Gumbel logits produced by prompt gate modules are used to determine whether to retrieve prompts from prompt pools. $L$ denoted each self-attention layer in the Transformer architectures. Outputs from the original CLIP model are represented in blue, while those incorporating retrieved prompts via the IA-GP strategy are shown in green. The processed visual and textual features are optimized through a contrastive loss function.
3 Approach
3.1 Preliminaries
Mutli-Domain Class-Incremental Learning. Given a pre-trained VLM, which incrementally learn from a stream of tasks originating from $\mathcal{T}$ distinct domains ${\{D_{1},D_{2},...,D_{\mathcal{T}}\}}$, each domain consists of ${N}$ samples, denoted as $D_{t}=(x^{t}_{n},y^{t}_{n})^{N}_{n=1}$, where ${x_{n}}$ is an image and $y_{n}$ is the corresponding one-hot ground truth. For domain $D_{n}$, the associated class set is given by $C_{n}=\{c^{n}_{i}\}^{M^{n}}_{i=1}$, where each $c^{n}_{i}$ represents a textual label. Under MCIL setting, the domain ${D_{n}}$ is accessible only during its corresponding $n$-th incremental learning phase. Moreover, class sets are disjoint across domains, i.e., $C_{i}\neq C_{j}$ for any $i\neq j$, and the data distributions also differ across domains, meaning $\mathbb{P}_{i}\neq\mathbb{P}_{j}$, where $\mathbb{P}$ denotes the distribution. During the test phase, the model is evaluated without accessing to the task identifier, requiring it to infer predictions across all previously seen domains without explicit domain information.
CLIP model. The pre-trained VLMs such as CLIP [33] consisted of an image encoder $F_{v}$ and a text encoder $F_{t}$. For a class $c_{i}^{n}$, CLIP model first transforms it to a sentence by a template such as “{a photo of {$c_{i}^{n}$}}”, and then encodes it into text embedding $t_{i}$. CLIP model is trained by leverages contrastive loss, the objective can be defined as:
$L=-\sum_{i=1}^{N}\log\left(\frac{\exp\left(\text{sim}\left(F_{v}(x_{i}),F_{t}(t_{i})\right)/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}\left(F_{v}(x_{i}),F_{t}(t_{j})\right)/\tau\right)}\right),$ (1)
where $\tau$ is the temperature, $\text{sim}(u,v)=\frac{u^{T}\cdot v}{\|u\|\|v\|}$ is the cosine similarity function, the contrastive loss enables the CLIP model to capture the inter-modal similarity between image embedding and all text embeddings.
Interference-free Knowledge Integration (IKI) mechanism. We follow  [36] to leverage a prompt-based incremental learning method. Specifically, a set of prompt pools are maintained for a stream of tasks, which can be denoted as:
$\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{\mathcal{T}}\},$ (2)
where $P_{t}=(K_{t},V_{t})$, and $K_{t},V_{t}\in\mathbb{R}^{l\times d}$. $l$ is the prompt length, $d$ is the embedding dimension. When a test sample $x_{n}$ comes, IKI mechanism first selects the corresponding prompt $\mathcal{P}_{r}$, and produces a task-specific attention output, which can be formulated as:
$O_{r}=\text{softmax}\left(\frac{Q_{n}K_{r}^{T}}{\sqrt{d}}\right)V_{r},$ (3)
where $O_{r},Q_{n}\in\mathbb{R}^{L\times d}$, and $Q_{n}$ is original query vector of CLIP model, $L$ is the length of embedded feature, $K_{r}$ and $V_{r}$ are come from the selected prompt $\mathcal{P}_{r}$. IKI mechanism leverages a residual branch for the top layers in a transformer architecture:
$O_{IKI}=O_{ori}+O_{r},$ (4)
where $O_{ori}$ is the original attention output of CLIP model. IKI mechanism incrementally learning different tasks by maintaining different prompt pools for each distribution.
3.2 Instance-Aware Prompting
3.2.1 Gated Prompting
In transformer architectures, the efficacy of a uniform prompting strategy across tasks with diverse distributions remains a critical challenge. For example, EuroSAT [14] (10 classes) and SUN397 [44] (397 classes) differ in distribution and granularity, and the differences become particularly significant when analyzing individual instances within a dataset, suggesting that prompting configurations must be adaptive to instance-specific characteristics. To address this, we propose Instance-Aware Gated Prompting (IA-GP), which employs an instance-aware gating mechanism to dynamically tailor the prompting strategy to the individual features.
The IA-GP strategy incorporates multiple prompting gates positioned before the vision transformer layers, as illustrated in Figure 2. Each prompt gate consists of a Gumbel linear function $H_{i}$, which maps the embedded image features to a $K$-dimensional space, and a Gumbel distribution which used to generate the samples uniformly. For each instance, we utilize the image features extracted by the original CLIP model as the input to the IA-GP module. This operation is motivated by the fact that the original CLIP model is frozen, ensuring that its raw image features reside in a stable distribution. We then compute the instance-aware Gumbel logit by:
$G_{i}=\frac{\exp(\log\left({H_{i}(F_{v}(x))+g_{i}})/{\tau}\right)}{\sum_{j=1}^{K}\exp(\log\left({H_{i}(F_{v}(x))+g_{j}})/{\tau}\right)}.$ (5)
In our approach, we set $K=2$ to implement a hard Gumbel Softmax mechanism, which facilitates binary decision-making for prompting. $g_{i}=-\log(-\log(U_{i}))$ represents a random logit sampled from the Gumbel distribution, where $U_{i}\sim U(0,1)$. The parameter $\tau$ denotes the temperature of the IA-GP module, controlling the sharpness of the decision boundary. The IA-GP module generates a two-dimensional output $G_{i}$, which is subsequently transformed into a one-hot vector via the hard Gumbel Softmax operation. This mechanism enables IAP approach to dynamically determine whether to prompt for each layer at the instance level. Specifically, if prompting is beneficial, the IA-GP module adjusts the Gumbel logit such that $G_{i}$ approaches 1. Conversely, if prompting operation is detrimental, the gate sets $G_{i}$ close to 0, thereby preserving the original output of the CLIP model. In summary, for each transformer layer during training, the output under our IA-GP strategy is formulated as follows:
$O_{IAP}=O_{ori}+\text{one-hot}(G)O_{r}.$ (6)
The operations described above are implemented for each individual instance. As a result, the IA-GP module can dynamically determining and assigning the appropriate the prompting layers according to the specific features of each instance.
3.2.2 Class-Distribution Driven Prompting
Incremental VLMs must mitigate backward forgetting (degradation of learned distributions) and prevent forward forgetting (impaired generalization to unseen distributions) to preserve zero-shot generalization. In the MCIL scenario, where task identifiers are absent during inference, distinguishing seen tasks from unseen tasks is crucial. DIKI [36] adjusts prompt weights using instance-to-task similarity as a confidence score. However, this approach is suboptimal, where high-confidence instances should use unadjusted prompts directly, while low-confidence instances risk noise if prompted, the original CLIP model needs to be employed. For instances with intermediate confidence scores, the confidence score needs to be more reliable. To address this, we propose Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP), a two-stage strategy that reassesses instances via both task- and class-specific perspectives during inference.
During the training phase, for each newly encountered task with the distribution $D_{i}$, the visual features of the images within $D_{i}$ are extracted using the original frozen CLIP image encoder. Subsequently, the mean vector $\mu_{i}$ and covariance matrix $\Sigma_{i}$ of these feature vectors are computed and stored:
$\mu_{i}=\mathbb{E}_{x_{i}\sim D_{i}}\left[F_{v}(x_{i})\right]$ (7)
$\Sigma_{i}=\mathbb{E}_{x_{i}\sim D_{i}}\left[\left(F_{v}(x_{i})-\mu_{i}\right)^{T}\left(F_{v}(x_{i})-\mu_{i}\right)\right],$
where $F_{v}(x_{i})$ represents the feature vector extracted for each image $x_{i}\in D_{i}$ by the frozen CLIP image encoder, ensuring that the features reside in a stable feature space due to the fixed encoder parameters. In the inference stage, for each learned distribution $D_{i}$, a multivariate Gaussian distribution $\mathcal{N}_{i}(\mu_{i},\Sigma_{i})$ is constructed using the stored $\mu_{i}$ and $\Sigma_{i}$. For a given test sample, its feature vector is extracted using the same CLIP image encoder, and the log-likelihood of this feature vector under each $\mathcal{N}_{i}$ is calculated to assess its fit to the learned distributions:
$\quad E_{i}^{\prime}=$ $\log\varphi(F_{v}(\hat{x});\mu_{i},\Sigma_{i})$ (8)
$\quad=-\frac{1}{2}\left[(F_{v}(\hat{x})-\mu_{i})^{T}\left(\Sigma_{i}\right)^{-1}(F_{v}(\hat{x})-\mu_{i})\right.$
$\quad+d\log 2\pi+\log|\Sigma_{i}|\Big{]},$
$\hat{x}\in\mathbb{R}^{d}$ denote the image input of a test sample, where $d$ represents the dimensionality of the image features. and $\varphi$ denote the probability density function (PDF) corresponding to the learned distribution associated with task $i$. The IKI mechanism computes the PDF values $E^{\prime}_{i}=\varphi(\hat{x})$ for each task $i$ and determines the maximum score as $E^{\prime}_{max}=\text{max}_{i\in[1,\mathcal{T}]}E^{\prime}_{i}$. Subsequently, it applies the sigmoid function $\sigma$ to get the confidence score, yielding $E_{max}=\sigma E^{\prime}_{max}$, which maps $E^{\prime}_{max}$ to the interval [0,1]. This value $E_{max}$ is then employed as the prompting weight within the attention layers of both the vision transformer and the text transformer. The IA-CDDP we proposed is designed with a two-stage architecture, as shown in Figure 5. In the first stage, the module evaluates the confidence scores of the test sample $\hat{x}$ across the $\mathcal{T}$ tasks. For samples exhibiting significantly high or low confidence scores, the module refrains from assigning a weight to the prompt. Instead, it directly utilizes either the incremental learning model or the original CLIP model, depending on the specific bounds:
Figure 3: Illustration of IA-CDDP strategy. The IA-CDDP strategy leverages visual features from the original frozen CLIP model for a two-stage confidence assessment per instance: 1) confidence scores between the instance and seen distributions are evaluated and binarized using predefined thresholds.2) for median confidence, the top $K$ highest scores are selected by modeling the in-task class distribution, with their mean computed as the final prompting weight.
    Method Aircraft [27] Caltech101 [10] CIFAR100 [22] DTD [6] EuroSAT [14] Flowers [28] Food [2] MNIST [8] OxfordPet [31] StanfordCars [21] SUN397 [44] Average
CLIP    Zero-shot 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
   Full Fine-tune 62.0 95.1 89.6 79.5 98.9 97.5 92.7 99.6 94.7 89.6 81.8 89.2
Transfer    Continual-FT 67.1 46.0 32.1 35.6 35.0 57.7 44.1 60.8 20.5 46.6 44.6
   LwF [26] 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 58.9
   iCaRL [34] 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
   LwF-VR [9] 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
   WiSE-FT [42] 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
   ZSCL [51] 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
   L2P [41] 65.6 50.9 30.4 41.4 49.3 71.8 36.3 77.5 55.3 53.4 53.2
   DualPrompt [40] 56.7 51.4 28.7 33.7 45.6 70.9 59.5 77.7 49.5 50.4 52.4
   S-Prompts [39] 67.3 49.4 26.7 39.7 47.1 70.2 34.3 78.9 56.7 52.2 52.2
   MoE-Adapter [47] 87.9 68.2 44.4 49.9 70.7 88.7 59.7 89.1 64.5 65.5 68.9
   DIKI [36] 92.9 69.1 43.2 43.9 65.4 85.3 56.0 88.4 64.0 65.6 67.4
   Ours 93.0 68.7 44.0 47.0 70.4 85.9 63.5 89.7 66.2 63.3 69.2
Average    Continual-FT 25.5 81.5 59.1 53.2 64.7 51.8 63.2 64.3 69.7 31.8 49.7 55.9
   LwF [26] 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
   iCaRL [34] 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
   LwF-VR [9] 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
   WiSE-FT [42] 26.7 86.5 64.3 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
   ZSCL [51] 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
   L2P [41] 38.0 85.2 78.2 61.3 72.9 74.9 79.7 59.1 82.0 59.7 55.4 67.9
   DualPrompt [40] 37.8 84.3 78.6 60.1 71.1 73.2 79.1 73.9 82.3 55.1 52.8 68.0
   S-Prompts [39] 37.5 92.5 77.5 58.2 76.4 74.1 78.8 57.9 83.0 60.8 54.4 68.3
   MoE-Adapter [47] 50.2 91.9 83.1 69.4 78.9 84.0 89.1 73.7 89.3 67.7 66.9 76.7
   DIKI [36] 45.4 95.7 83.0 65.0 78.2 82.5 87.1 71.7 90.0 67.2 66.6 75.7
   Ours 45.9 95.8 83.3 66.5 79.5 84.8 87.5 76.6 91.0 69.2 64.5 76.8
Last    Continual-FT 31.0 89.3 65.8 67.3 88.9 71.1 85.6 99.6 92.9 77.3 81.1 77.3
   LwF [26] 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
   iCaRL [34] 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
   LwF-VR [9] 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
   WiSE-FT [42] 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
   ZSCL [51] 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
   L2P [41] 38.0 87.1 84.2 72.9 86.0 96.1 89.2 99.0 94.1 79.6 76.0 82.0
   DualPrompt [40] 37.8 87.1 84.6 71.8 89.2 96.3 89.1 99.1 94.5 79.9 76.5 82.3
   S-Prompts [39] 37.5 95.1 83.7 70.2 97.5 96.5 89.0 99.1 94.0 79.5 75.8 83.4
   MoE-Adapter[47] 49.8 92.2 86.1 78.1 95.7 94.3 89.5 98.1 89.9 81.6 80.0 85.0
   DIKI [36] 45.4 95.9 86.0 73.0 97.8 96.8 89.3 99.3 94.4 81.8 76.4 85.1
   Ours 46.8 96.1 86.7 75.2 98.1 97.0 89.6 99.4 94.7 82.8 76.7 85.7
Table 1: Comparison with SOTA on MCIL benchmark in terms of “Transfer”, “Average”, and “Last” metrics (%). “Ours” denotes our method. We label the best and second methods with bold and underline styles. The presented results are derived from the Order-I, for Order-II results, please refer to the supplementary materials.
$\hat{E}_{max}=\begin{cases}0&\text{for }E_{max}\leq\text{lower bound}\\ 1&\text{for }E_{max}\geq\text{upper bound}\\ E_{max}&\text{into the second stage}\end{cases}\quad\quad.$ (9)
For samples that yield confidence scores within a median range, we employ the second stage of our IA-CDDP approach, which reconstructs the Gaussian distribution for each class within the distribution $D_{i}$:
$\quad E^{\prime}_{i,j}=$ $\log\varphi(F_{v}(\hat{x});\mu_{i,j},\Sigma_{i,j})$ (10)
$\quad=-\frac{1}{2}\left[(F_{v}(\hat{x})-\mu_{i,j})^{T}\left(\Sigma_{i,j}\right)^{-1}(F_{v}(\hat{x})-\mu_{i,j})\right.$
$\quad+d\log 2\pi+\log|\Sigma_{i,j}|\Big{]},$
where $\mu_{i,j}$ and $\Sigma_{i,j}$ denote the mean and covariance matrix of class $j$ within distribution $D_{i}$. For every specific instance, We then compute the mean of the confidence scores associated with the most similar $K$ classes. The mean of these $K$ classes serves as the final weight for the prompt, which is subsequently integrated into the attention layers of the transformer model:
$\hat{E}_{means}=\frac{1}{K}\sum_{k=1}^{K}\sigma E^{\prime}_{[i]},$ (11)
where $E_{[i]}$ represent the top $K$ confidence scores corresponding to the most similar classes within the distribution. Through the application of the second stage of the IA-CDDP strategy, test samples that are distant from the Gaussian distribution of the entire task can still obtain a reasonable confidence score.
Furthermore, for prompts in text encoder, we follow  [36] and leverage a batch-wise confidence score to align with the vision encoder, which can be denoted as follow:
${E}_{txt}=\frac{1}{B}\sum_{b=1}^{B}\sigma E^{\prime}_{b},$ (12)
where $B$ is the batchsize, and $E^{\prime}_{b}$ is the confidence score of images in a batch.
By integrating the IA-GP strategy with IA-CDDP, VLMs can incrementally acquire knowledge from a stream of tasks. When the distribution of a task is identified as ID, the model assigns prompts to the relevant layers and applies the corresponding weights. This approach mitigates backward forgetting of previously encountered distributions. Conversely, when the distribution is classified as OOD, the model refrains from utilizing prompts, thereby preserving the initial output of the model. This leverages the zero-shot generalization ability of the pre-trained model, effectively reducing forward forgetting for unseen tasks.
4 Experiments
4.1 Experimental Setting
Dataset. We evaluate our approach within the context of a Multi-Domain Class-Incremental Learning benchmark comprising 11 diverse datasets: Aircraft [27], Caltech101 [10], CIFAR100 [22], DTD [6], EuroSAT [14], Flowers [28], Food [2], MNIST [8], OxfordPet [31], StanfordCars [21] and SUN397 [44]. These datasets collectively encompass 1201 classes, each characterized by distinct distributions. We follow  [51] and conduct our experiment through Order-I and Order-II. Experiments are done in Order-I by default while the experimental details of Order-II is available in the supplementary materials.
Figure 4: Mean prompting layers for different tasks. The tasks with complex distributions tend to allocate more prompting layers.
Metrics. We evaluate our method’s performance using three metrics proposed by  [51]: “Transfer”, “Last”, and “Avg”. These metrics are tailored to assess distinct aspects of incremental learning, collectively offering a robust evaluation framework. The details about these metrics are included in the supplementary material.
Comparison methods. We compare our proposed IAP method with two categories of SOTA methods in incremental learning. The first category encompasses full-parameter fine-tuning methods, which typically leverage knowledge distillation or rehearsal-based techniques and update all parameters of the model. Notable approaches in this category include Continual-FT, LwF [26], iCaRL [34], LwF-VR [9], WiSE-FT [42], and ZSCL [51]. The second category comprises PEFT incremental learning methods, which adapt models to new distributions by maintaining only a small set of additional parameters. Prominent methods in this category include L2P [41], DualPrompt [40], S-Prompt [39], MoE-Adapter [47] and DIKI [36]. Our proposed IAP method also falls within the PEFT incremental learning benchmark.
Implementation details. To ensure equitable comparisons, we maintain the same backbone as  [51], utilizing CLIP ViT-B/16 as our vision-language model. We employ Stochastic Gradient Descent (SGD) as the optimizer with an initial learning rate of 5.0, paired with a cosine annealing scheduler to dynamically adjust the learning rate during training. The model is trained for 10 epochs across all tasks. Following  [36], we apply prompts to the first 8 transformer layers of the text encoder, with a prompt length of 8. For the image encoder, we adopt the IA-GP strategy and integrate the prompts into all transformer layers during training, with the Gumbel sampling temperature set to 3.0. In the first stage of the IA-CDDP module, we employ an upper bound of 0.8 and a lower bound of 0.2. In the second stage, we set $K=5$ to select the top five most relevant classes. All experiments are performed on an NVIDIA 4090 GPU.
4.2 Experimental Results
Main results. The main results are presented in Table 1. We provide a detailed performance comparison of our proposed method against other SOTA methods within the MCIL benchmark. At the top of the table, we report the zero-shot generalization ability of the CLIP model and its performance under full-parameter fine-tuning as baselines. Our proposed method, denoted as “Ours” in the table, outperforms all SOTA methods for MCIL across three key metrics. Specifically, our method achieves the highest performance on the “Transfer” metric for 4 out of 11 tasks, the “Average” metric for 7 tasks, and the “Last” metric for 6 tasks. Furthermore, our approach establishes new SOTA performance in terms of the overall mean values across all three metrics. While the MoE-Adapter [47] method demonstrates the closest performance to ours, however, our approach achieves this with a significantly smaller number of learnable parameters. For a detailed analysis of the experimental results, please refer to the subsequent sections.
Method Memory Parameters GPU
LwF [26] 211M 32172 MiB
LwF-VR [9] 211M 32236 MiB
ZSCL [51] 211M 26290 MiB
MoE-Adapter [47] 59.8 M 22358 MiB
Ours 2.4 M 19610 MiB
Table 2: Comparison of the use of memory buffers, the scale of trainable parameters, and the GPU burden during the training process.
Method Transfer Average Last
DIKI 67.4 75.7 85.1
DIKI* 68.7 76.3 85.1
IAP w/o IA-CDDP 68.6 76.6 85.5
IAP w/o IA-GP 69.1 76.5 85.3
IAP 69.2 76.8 85.7
Table 3: The ablation experiments for each module of our proposed method are presented below. Asterisk (*) denotes the original performance reported in the DIKI paper.
Figure 5: The process of our IA-CDDP module for test samples across tasks 1, 4, and 9. As illustrated in the first row of the figure, the first stage of IA-CDDP constructs distributions at the task level and performs inference based on the similarity between samples and each distribution. Test samples with either excessively low or high confidence scores are directly classified, while samples with intermediate confidence scores proceed to the second stage. At this stage, we construct distributions for all classes within the task and calculates similarities with the 5 nearest class distributions, highlighted by the red dashed circle in the second row. The mean of these distribution similarities serves as the confidence score for the test samples, which subsequently guides the following prompting process.
Trainable parameters and memory buffer. We present a comparison of the trainable parameters and rehearsal memory buffer requirements for our method and four other SOTA approaches in Table 2. Leveraging PEFT strategies, our method eliminates the need for external storage of representative samples or features. Furthermore, our approach utilizes only $1/25$ of the trainable parameters compared to the MoE-Adapter [47] method. This substantial reduction is achieved because our method does not require the training of numerous experts; instead, for each task, only the corresponding prompt pool and prompt gate module are trained.
Effect of Instance-Aware Gated Prompting strategy. We conducted a statistical analysis of mean prompting layers employed by our IA-GP strategy across different distributions, as illustrated in Figure 4. The red line represents the prompt addition strategy employed by DIKI [36], which consistently prompting the first eight transformer layers. In contrast, our IAP method, benefiting from the IA-GP strategy, demonstrates greater flexibility by dynamically determining whether to apply prompts to different transformer layers at the instance level. Experimental results indicate that for tasks with more complex distributions, such as SUN397 (397 classes) and Food101 (101 classes), the IA-GP strategy tends to incorporate more prompts to effectively capture intricate distributional information. Conversely, for relatively simpler datasets like DTD (47 classes) and EuroSAT (10 classes), the strategy selectively employs fewer prompts, thereby achieving superior performance while reducing inference time. We also provide the specific prompting layers for each task, presented as a heatmap. For details, please refer to the supplementary materials.
Effect of Instance-Aware Class-Distribution-Driven Prompting strategy. In Figure 5, we visualize the distributions of all 11 tasks and the class distributions of three specific tasks: Aircraft (100 classes), OxfordPet (47 classes), and StanfordCars (196 classes). For each task, we use a corresponding instance as an example, denoted by a star. The reference distance, displayed in the lower left corner of each graph, serves as an indicator of confidence score, with values closer to 1 signifying higher confident. The top row of the figure illustrates the distributions of the 11 tasks. The DIKI [36] method leverages the similarity between an instance and the tasks distribution to derive confidence scores, which in turn control the weight of subsequent prompting operation. In contrast, our IA-CDDP method incorporates a second stage that computes the average of the five nearest class distributions (highlighted by red dotted circles) at the instance level. The reference distances demonstrate that our IA-CDDP method assigns more appropriate instance-aware confidence scores, thereby enhancing the efficacy of subsequent prompts.
Quantitative Analysis Our method is implemented on a classic prompt-based MCIL method DIKI [36]  [^1] https://github.com/lloongx/DIKI. We perform a quantitative analysis of our proposed IA-GP and IA-CDDP modules and the results are shown in Table 3. For the baseline method, we observed that the reproduced results (DIKI) are lower than the papers’ results (DIKI*). Compared to baseline DIKI, both IA-GP and IA-CDDP can achieve a performance improvement. The details are included in the supplemental material.
5 Conclusion
In this paper, we propose the Instance-Aware Prompting method to tackle the challenges of backward forgetting and forward forgetting faced by pre-trained VLMs during incremental learning of new tasks. The IAP method comprises two meticulous strategies designed at the instance level. The Instance-Aware Gated Prompting strategy dynamically controls whether to prompt at the instance level by incorporating a prompt gate module for each transformer layer. The Instance-Aware Class-Distribution-Driven Prompting strategy assigns more precise weights to each instance through a two-stage determination process. By integrating these two strategies, our method simultaneously resolves the questions of whether to apply prompt and how much prompt to apply, enabling pre-trained VLMs to adapt more flexibly when incrementally learning new tasks. Extensive experimental results demonstrate that the IAP method represents the current SOTA approach in the MCIL benchmark.